• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name dwang

    Log entry time 00:19:41 on October 2,2009

    Entry number 292668

    keyword=compton Edet current at 1 mm

    The electron detector is hot at 0mm, so we decided to move it to 1mm for now. May need more tune later in order to get closer to the beam.

    FIGURE 1